# Inferring Change Points in High-Dimensional Regression via Approximate Message Passing

### v2

@@ C.2.2 Computation of $g^{*t}$ for Logistic Regression
$\mu_{\ell}\equiv\mu_{\ell}(\bm{V}_{i}):=\left({\mathop{\mathbb{E}}}[\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}=\bm{V}_{i}]\right)_{\ell}\,,\qquad\sigma_{\ell}\equiv\sigma_{\ell}(\bm{V}_{i}):=\sqrt{\left[{\rm{{Cov}}}(\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}=\bm{V}_{i})\right]_{\ell,\ell}}\,.$ (90)
@@
Lemma 12 (Approximation of logistic function, Section 4.5.2 in Bishop (2006))
@@ C.2.3 Computation of $g^{*t}$ for Rectified Linear Regression
In the rest of this subsection, we provide the expression for ${\mathcal{L}}_{1}(\bm{V}_{i},u_{i},\omega|\ell)$ and $\widetilde{\lambda}^{t}(k,\bm{V}_{i},u_{i},\ell,\omega)$ for $\omega=0$ or 1, respectively.
@@
$=\phi\left(\bm{V}_{i};{\bm{0}},{\bm{\Sigma}}_{\bm{V}}^{t}\right)\phi(u_{i};\mu_{\ell},\sigma_{\ell}^{2}+\sigma^{2})\Phi(\mu_{\ell*}/\sigma_{\ell*})\left\{\frac{\lambda_{k\ell}}{\sigma^{2}_{\ell}}\left[\mu_{\ell*}+\sigma_{\ell*}\frac{\phi(0;\mu_{\ell*},\sigma^{2}_{\ell*})}{\Phi(\mu_{\ell*}/\sigma_{\ell*})}\right]+\mu_{k}-\frac{\lambda_{k\ell}}{\sigma^{2}_{\ell}}\mu_{\ell}\right\}.$ (109)
@@ D.1 Proof of Proposition 3
Proof  Recall that a signal configuration vector ${\bm{\Psi}}\in\mathcal{X}$ is a vector that is piece-wise constant with respect to its indices $i\in[n]$, with jumps of size $1$ occurring at the indices $\{\eta_{\ell}\}_{\ell=1}^{L^{*}-1}$. Without loss of generality, we assume ${\bm{\Psi}}$ is also monotone (otherwise, any non-distinct signal can be treated as a new signal having perfect correlation with the first). Recall that ${\bm{\eta}}\in[n]^{L^{*}-1}$ is the change point vector corresponding to ${\bm{\Psi}}$, i.e., ${\bm{\eta}}=U^{-1}({\bm{\Psi}})$, and that for $i\in[n]$ the distance of $\eta_{i}$ to a change point estimate $\hat{{\bm{\eta}}}\in[n]^{L-1}$ is defined as $d(\eta_{i},\{\hat{\eta}_{j}\}_{j=1}^{L-1}):=\min_{\hat{\eta}_{j}\in\{\hat{\eta}_{j}\}_{j=1}^{L-1}}\|\eta_{i}-\hat{\eta}_{j}\|_{2}$. We then have that $d(\eta_{i},\{\hat{\eta}_{j}\}_{j=1}^{L-1})\leq\|U({\bm{\eta}})-U(\hat{{\bm{\eta}}})\|^{2}_{F}$ for all $i\in[L^{*}-1]$, and similarly $d(\{\eta_{i}\}_{i=1}^{L^{*}-1},\hat{\eta}_{j})\leq\|U({\bm{\eta}})-U(\hat{{\bm{\eta}}})\|^{2}_{F}$ for all $j\in[L-1]$. This implies that $d_{H}({\bm{\eta}},\hat{{\bm{\eta}}})\leq\|U({\bm{\eta}})-U(\hat{{\bm{\eta}}})\|_{F}^{2}$.
We first prove that $(\bm{\Theta}^{t},\bm{X}\bm{B})\mapsto d_{H}({\bm{\eta}},\hat{{\bm{\eta}}}(\bm{\Theta}^{t},q(\bm{X}\bm{B},{\bm{\Psi}},\bm{\varepsilon}))$ is uniformly pseudo-Lipschitz. Consider two sets of inputs, $\bm{A}^{(1)}:=\begin{bmatrix}(\bm{\Theta}^{t})^{(1)}&(\bm{X}\bm{B})^{(1)}\end{bmatrix}\in{\mathbb{R}}^{n\times 2L}$ and $\bm{A}^{(2)}:=\begin{bmatrix}(\bm{\Theta}^{t})^{(2)}&(\bm{X}\bm{B})^{(2)})\end{bmatrix}\in{\mathbb{R}}^{n\times 2L}$, and let $\hat{{\bm{\eta}}}(\bm{A}^{(1)}):=\hat{{\bm{\eta}}}\left((\bm{\Theta}^{t})^{(1)},q((\bm{X}\bm{B})^{(1)},{\bm{\Psi}},\bm{\varepsilon})\right)$, $\hat{{\bm{\eta}}}(\bm{A}^{(2)}):=\hat{{\bm{\eta}}}\left((\bm{\Theta}^{t})^{(2)},q((\bm{X}\bm{B})^{(2)},{\bm{\Psi}},\bm{\varepsilon})\right)$. We then have that:
@@ Appendix E Further Implementation and Experiment Details
where $\hat{{\mathop{\mathbb{E}}}}$ denotes an expectation estimate via Monte Carlo. For example, in the case of (115), we generate 300 to 1000 independent samples of $(\bm{Z}_{1},(\bm{G}_{\bm{\Theta}}^{t})_{1},\bar{\varepsilon})$, with ${\bm{Z}}_{1}\sim{\sf N}({\bm{0}},{\bm{\rho}})$, $(\bm{G}_{\bm{\Theta}}^{t})_{1}\mathbin{\overset{}{\kern 0.0pt\sim}}{\sf N}({\bm{0}},{\bm{\kappa}}_{\bm{\Theta}}^{t,t})$, and $\bar{\varepsilon}\sim{\mathbb{P}}_{\bar{\varepsilon}}$. We form $(\bm{V}_{\bm{\Theta}}^{t})_{1}$ according to the first row of (7), i.e., $(\bm{V}_{\bm{\Theta}}^{t})_{1}=\bm{Z}_{1}{\bm{\rho}}^{-1}{\bm{\nu}}_{\bm{\Theta}}^{t}+(\bm{G}_{\bm{\Theta}}^{t})_{1}$. For $\ell\in\{0,\dots,L-1\}$, this yields a set of samples of the random variables $g_{i}^{t}\left((\bm{V}_{\bm{\Theta}}^{t})_{1},q_{1}(\bm{Z}_{1},\Psi_{\eta_{\ell}},\bar{\varepsilon})\right)$ and $\partial_{1}\tilde{g}_{i}^{t}(\bm{Z}_{1},(\bm{V}_{\bm{\Theta}}^{t})_{1},\Psi_{\eta_{\ell}},\bar{\varepsilon})$ (the latter function is computed using Automatic Differentiation (AD)). We then compute $\hat{{\mathop{\mathbb{E}}}}$ in (114) and (115) by averaging. The $\hat{{\mathop{\mathbb{E}}}}$ terms in (116)–(117) are similarly computed.
@@
Figure 8: Comparison between AMP and $\mathtt{DCDP}$ in the experimental setting of Figure 3(a). Solid red: $\mathtt{DCDP}$ with hyperparameters chosen using CV. Solid green: AMP using $L=3$, the optimal denoiser $f^{t}$ with the sparsity level estimated using CV. Solid black: AMP using the true signal prior.
@@
Figure 8 shows results from an additional set of experiments comparing AMP with $\mathtt{DCDP}$ (the fastest algorithm in Figure 3(a)) for the linear model, using different change point priors $\pi_{\bar{{\bm{\Psi}}}}$ (Figure 8(a)) or denoisers $f^{t}$ (Figure 8(b)). In both figures, the solid red plot shows $\mathtt{DCDP}$ performance with hyperparameters chosen using CV. Solid black corresponds to AMP using the true signal prior. Solid green corresponds to AMP using $L=3$, the optimal denoiser $f^{t}$ with the sparsity level estimated using CV. In Figure 8(a), AMP performs slightly worse with $L=4$ instead of $L=3$, because the prior assigns non-zero probability to the change point configurations with $L=4$ signals, which is mismatched from the ground truth $L^{*}=3$. In Figure 8(b), AMP performs slightly worse at lower $\delta$ when using a suboptimal soft thresholding (ST) denoiser. Nevertheless, in both Figures 8(a) and 8(b), AMP largely outperforms $\mathtt{DCDP}$ despite the suboptimal choices of prior or denoiser.
Figure 9 compares AMP with $\mathtt{DCDP}$ for $n=500$ and varying $p$, both using hyperparameters chosen via cross-validation. The runtime shown is the average runtime per set of CV parameters. For a fixed set of hyperparameters, the time complexity of $\mathtt{DCDP}$ scales as LASSO$(n,p)$ compared to $O(np)$ for AMP. The runtime of $\mathtt{DCDP}$ therefore grows faster with increasing $p$.
@@
For synthetic data experiments in Section 4, we initialize AMP with $f^{0}(\bm{B}^{0})$ sampled row-wise independently from the prior ${\mathbb{P}}_{\bar{\bm{B}}}$ used to define the ensemble state evolution (20)–(23). Figures 3(a), 4, 6, 8 and 9 use Bernoulli-Gaussian priors taking the form ${\mathbb{P}}_{\bar{\bm{B}}}=\alpha{\sf N}({\bm{0}},\sigma^{2}{\bm{I}})+(1-\alpha)\delta_{{\bm{0}}}$ with $\alpha\in(0,1)$. Figure 4 uses $\alpha=1/6$ and $\sigma=2.5$. Figures 3(a), 6, 8 and 9 use $\alpha=0.5$ and $\sigma^{2}=\delta$ to ensure that the signal power ${\mathop{\mathbb{E}}}[(\bm{X}_{i})^{\top}{\bm{\beta}}^{(i)}]^{2}$ is held constant for varying $\delta$, and they use two change points at $n/3$ and $158n/$ 15 as the ground truth. In Figures 3(a), 8 and 9, AMP uses cross validation (CV) over 5 values of $\hat{\alpha}:\{0.1,0.3,0.45,0.6,0.9\}$ which do not contain the true $\alpha=0.5$. $\mathtt{DCDP}$, $\mathtt{DPDU}$, and $\mathtt{DP}$ each have two hyperparameters: one corresponding to the $\ell_{1}$ penalty and the other penalizing the number of change points. We run cross-validation on these hyperparameters for $12$, $12$, or $42$ pairs of values, respectively, using the suggested values from the original papers.
For the experiments in Figure 3(b), we consider signals with sparse changes between adjacent signals. The prior takes the following form:
[…]
@@
[…]
@@ Appendix F Computational Cost
For $i\in[n]$, the function $g_{i}^{*t}$ in Proposition 2 only depends on $\bar{{\bm{\Psi}}}$ in (85) through the marginal probability of $\bar{\Psi}_{i}$, i.e., $\pi_{\bar{\Psi}_{i}}:\ell\mapsto\sum_{\{{\bm{\psi}}:\psi_{i}=\ell\}}\pi_{\bar{{\bm{\Psi}}}}({\bm{\psi}})$. This means $g_{i}^{*t}$ can be efficiently computed, involving only a sum over $\bar{\Psi}_{i}\in[L]$. Indeed, from the implementation details in Appendix C.2, both $f_{j}^{*t}$ and $g_{i}^{*t}$ can be computed in $O(L^{3})$ time for each $i,j$. Thus $f^{*t}$ and $g^{*t}$ can be computed in $O(nL^{3})$. The per-iteration computational cost of AMP is therefore dominated by the matrix multiplications in (LABEL:eq:amp), which are $O(npL)$.
@@
For ${\bm{\Psi}}\in\mathcal{X}$, the denominator in (33) only requires evaluating a polynomial number of terms, (${n\choose{L-1}}$. Therefore, choosing $f^{t},g^{t}$ as in Proposition 2, the computational complexity of computing $p_{\bar{{\bm{\Psi}}}|\bar{\bm{V}}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}({\bm{\psi}}|\bm{\Theta}^{t},\bm{y})$ from (35) along with AMP is $O(npL+nL^{3}+n^{L-1})=O(np+n^{L-1})$. Further, assuming the expectations in (10)–(13) can be well-approximated in $O(n)$ time, $p_{\bar{{\bm{\Psi}}}|{\bm{V}}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}({\bm{\psi}}|\bm{\Theta}^{t},\bm{y})$ from (33) along with AMP can be computed in $O(np+n^{L-1})$ time.

### v3

@@ C.2.2 Computation of $g^{*t}$ for Logistic Regression
$\mu_{\ell}\equiv\mu_{\ell}(\bm{V}_{i}):=\left({\mathop{\mathbb{E}}}[\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}=\bm{V}_{i}]\right)_{\ell}\,,\quad\sigma_{\ell}\equiv\sigma_{\ell}(\bm{V}_{i}):=\sqrt{\left[{\rm{{Cov}}}(\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}=\bm{V}_{i})\right]_{\ell,\ell}}\,.$ (93)
@@
Lemma 12 (Logistic approximation, Section 4.5.2 in Bishop (2006))
@@ C.2.3 Computation of $g^{*t}$ for Rectified Linear Regression
In the rest of this subsection, we provide the expression for the terms ${\mathcal{L}}_{1}(\bm{V}_{i},u_{i},\omega|\ell)$ and $\widetilde{\lambda}^{t}(k,\bm{V}_{i},u_{i},\ell,\omega)$ for $\omega=0$ or 1, respectively.
@@
$=\phi\left(\bm{V}_{i};{\bm{0}},{\bm{\Sigma}}_{\bm{V}}^{t}\right)\phi(u_{i};\mu_{\ell},\sigma_{\ell}^{2}+\sigma^{2})\Phi(\mu_{\ell*}/\sigma_{\ell*})$
$\hskip 170.71652pt\cdot\left\{\frac{\lambda_{k\ell}}{\sigma^{2}_{\ell}}\left[\mu_{\ell*}+\sigma_{\ell*}\frac{\phi(0;\mu_{\ell*},\sigma^{2}_{\ell*})}{\Phi(\mu_{\ell*}/\sigma_{\ell*})}\right]+\mu_{k}-\frac{\lambda_{k\ell}}{\sigma^{2}_{\ell}}\mu_{\ell}\right\}.$ (112)
@@ D.1 Proof of Proposition 3
Proof  Recall that a signal configuration vector ${\bm{\Psi}}\in\mathcal{X}$ is a vector that is piece-wise constant with respect to its indices $i\in[n]$, with jumps of size $1$ occurring at the indices $\{\eta_{\ell}\}_{\ell=1}^{L^{*}-1}$. Without loss of generality, we assume $\bm{\Psi}$} is also monotone (otherwise, any non-distinct signal can be treated as a new signal having perfect correlation with the first). Recall that ${\bm{\eta}}\in[n]^{L^{*}-1}$ is the change point vector corresponding to $\bm{\Psi}$}, that is, ${\bm{\eta}}=U^{-1}({\bm{\Psi}})$, and that for $i\in[n]$ the distance of $\eta_{i}$ to a change point estimate $\hat{{\bm{\eta}}}\in[n]^{L-1}$ is defined as $d(\eta_{i},\{\hat{\eta}_{j}\}_{j=1}^{L-1}):=\min_{\hat{\eta}_{j}\in\{\hat{\eta}_{j}\}_{j=1}^{L-1}}\|\eta_{i}-\hat{\eta}_{j}\|_{2}$. We then have that $d(\eta_{i},\{\hat{\eta}_{j}\}_{j=1}^{L-1})\leq\|U({\bm{\eta}})-U(\hat{{\bm{\eta}}})\|^{2}_{F}$ for all $i\in[L^{*}-1]$, and similarly $d(\{\eta_{i}\}_{i=1}^{L^{*}-1},\hat{\eta}_{j})\leq\|U({\bm{\eta}})-U(\hat{{\bm{\eta}}})\|^{2}_{F}$ for all $j\in[L-1]$. This implies that $d_{H}({\bm{\eta}},\hat{{\bm{\eta}}})\leq\|U({\bm{\eta}})-U(\hat{{\bm{\eta}}})\|_{F}^{2}$.
We first prove that $(\bm{\Theta}^{t},\bm{X}\bm{B})\mapsto d_{H}({\bm{\eta}},\hat{{\bm{\eta}}}(\bm{\Theta}^{t},q(\bm{X}\bm{B},{\bm{\Psi}},\bm{\varepsilon}))/n$ is uniformly pseudo-Lipschitz. Consider two sets of inputs, $\bm{A}^{(1)}:=\begin{bmatrix}(\bm{\Theta}^{t})^{(1)}&(\bm{X}\bm{B})^{(1)}\end{bmatrix}\in{\mathbb{R}}^{n\times 2L}$ and $\bm{A}^{(2)}:=\begin{bmatrix}(\bm{\Theta}^{t})^{(2)}&(\bm{X}\bm{B})^{(2)})\end{bmatrix}\in{\mathbb{R}}^{n\times 2L}$, and let $\hat{{\bm{\eta}}}(\bm{A}^{(1)}):=\hat{{\bm{\eta}}}\left((\bm{\Theta}^{t})^{(1)},q((\bm{X}\bm{B})^{(1)},{\bm{\Psi}},\bm{\varepsilon})\right)$, $\hat{{\bm{\eta}}}(\bm{A}^{(2)}):=\hat{{\bm{\eta}}}\left((\bm{\Theta}^{t})^{(2)},q((\bm{X}\bm{B})^{(2)},{\bm{\Psi}},\bm{\varepsilon})\right)$. We then have that:
@@ Appendix E Further Implementation and Experiment Details
where $\hat{{\mathop{\mathbb{E}}}}$ denotes an expectation estimate via Monte Carlo. For example, in the case of (118), we generate 300 to 1000 independent samples of $(\bm{Z}_{1},(\bm{G}_{\bm{\Theta}}^{t})_{1},\bar{\varepsilon})$, with ${\bm{Z}}_{1}\sim{\sf N}({\bm{0}},{\bm{\rho}})$, $(\bm{G}_{\bm{\Theta}}^{t})_{1}\mathbin{\overset{}{\kern 0.0pt\sim}}{\sf N}({\bm{0}},{\bm{\kappa}}_{\bm{\Theta}}^{t,t})$, and $\bar{\varepsilon}\sim{\mathbb{P}}_{\bar{\varepsilon}}$. We form $(\bm{V}_{\bm{\Theta}}^{t})_{1}$ according to the first row of (7), that is, $(\bm{V}_{\bm{\Theta}}^{t})_{1}=\bm{Z}_{1}{\bm{\rho}}^{-1}{\bm{\nu}}_{\bm{\Theta}}^{t}+(\bm{G}_{\bm{\Theta}}^{t})_{1}$. For $\ell\in\{0,\dots,L-1\}$, this yields a set of samples of the random variables $g_{i}^{t}\left((\bm{V}_{\bm{\Theta}}^{t})_{1},q_{1}(\bm{Z}_{1},\Psi_{\eta_{\ell}},\bar{\varepsilon})\right)$ and $\partial_{1}\tilde{g}_{i}^{t}(\bm{Z}_{1},(\bm{V}_{\bm{\Theta}}^{t})_{1},\Psi_{\eta_{\ell}},\bar{\varepsilon})$ (the latter function is computed using Automatic Differentiation (AD)). We then compute $\hat{{\mathop{\mathbb{E}}}}$ in (117) and (118) by averaging. The $\hat{{\mathop{\mathbb{E}}}}$ terms in (119)–(120) are similarly computed.
@@
(a) Solid purple: AMP using $L=3,4$ and the optimal denoiser $f^{t}$ with the sparsity level estimated using CV.
(b) Dashed purple: AMP using $L=3$ and a suboptimal soft thresholding (ST) denoiser $f^{t}$ whose threshold is selected using CV.
Figure 8: Comparison between AMP and $\mathtt{DCDP}$ in the experimental setting of Figure 3a. Solid red: $\mathtt{DCDP}$ with hyperparameters chosen using CV. Solid green: AMP using $L=3$, the optimal denoiser $f^{t}$ with the sparsity level estimated using CV. Solid black: AMP using the true signal prior.
@@
Figure 8 shows results from an additional set of experiments comparing AMP with $\mathtt{DCDP}$ (the fastest algorithm in Figure 3a) for the linear model, using different change point priors $\pi_{\bar{{\bm{\Psi}}}}$ (Figure 8a) or denoisers $f^{t}$ (Figure 8b). In both figures, the solid red plot shows $\mathtt{DCDP}$ performance with hyperparameters chosen using CV. Solid black corresponds to AMP using the true signal prior. Solid green corresponds to AMP using $L=3$, the optimal denoiser $f^{t}$ with the sparsity level estimated using CV. In Figure 8a, AMP performs slightly worse with $L=4$ instead of $L=3$, because the prior assigns non-zero probability to the change point configurations with $L=4$ signals, which is mismatched from the ground truth $L^{*}=3$. In Figure 8b, AMP performs slightly worse at lower $\delta$ when using a suboptimal soft thresholding (ST) denoiser. Nevertheless, in both Figures 8a and 8b, AMP largely outperforms $\mathtt{DCDP}$ despite the suboptimal choices of prior or denoiser.
Figure 9 compares AMP with $\mathtt{DCDP}$ for $n=500$ and varying $p$, both using hyperparameters chosen via cross-validation. The runtime shown is the average runtime per set of CV parameters. For a fixed set of hyperparameters, the time complexity of $\mathtt{DCDP}$ scales as LASSO(n,p)(n,p) compared to $O(np)$ for AMP. The runtime of $\mathtt{DCDP}$ therefore grows faster with increasing $p$.
@@
For synthetic data experiments in Section 4, we initialize AMP with $f^{0}(\bm{B}^{0})$ sampled row-wise independently from the prior ${\mathbb{P}}_{\bar{\bm{B}}}$ used to define the ensemble state evolution (20)–(23). Figures 3a, 4, 6, 8 and 9 use Bernoulli-Gaussian priors taking the form ${\mathbb{P}}_{\bar{\bm{B}}}=\alpha{\sf N}({\bm{0}},\sigma^{2}{\bm{I}})+(1-\alpha)\delta_{{\bm{0}}}$ with $\alpha\in(0,1)$. Figure 4 uses $\alpha=1/6$ and $\sigma=2.5$. Figures 3a, 6, 8 and 9 use $\alpha=0.5$ and $\sigma^{2}=\delta$ to ensure that the signal power ${\mathop{\mathbb{E}}}[(\bm{X}_{i})^{\top}{\bm{\beta}}^{(i)}]^{2}$ is held constant for varying $\delta$, and they use two change points at $n/3$ and $8n/15$ as the ground truth. In Figures 3a, 8 and 9, AMP uses cross validation (CV) over 5 values of $\hat{\alpha}:\{0.1,0.3,0.45,0.6,0.9\}$ which do not contain the true $\alpha=0.5$. $\mathtt{DCDP}$, $\mathtt{DPDU}$, and $\mathtt{DP}$ each have two hyperparameters: one corresponding to the $\ell_{1}$ penalty and the other penalizing the number of change points. We run cross-validation on these hyperparameters for $12$, $12$, or $42$ pairs of values, respectively, using the suggested values from the original papers.
For the experiments in Figure 3b, we consider signals with sparse changes between adjacent signals. The prior takes the following form:
$\beta^{(\eta_{0})}_{j}\mathbin{\overset{\text{i.i.d.}}{\kern 0.0pt\sim}}{\sf N}(0,\kappa^{2})$
@@
where $w_{\ell}\mathbin{\overset{\text{i.i.d.}}{\kern 0.0pt\sim}}{\sf N}(0,\sigma_{w}^{2})$ creates the sparse change between adjacent signals, and $\nu:=\sqrt{\frac{\kappa^{2}}{\kappa^{2}+\sigma_{w}^{2}}}$ is a rescaling factor that ensures uniform signal magnitude ${\mathop{\mathbb{E}}}[(\beta_{j}^{(\eta_{0})})^{2}]=\dots={\mathop{\mathbb{E}}}[(\beta_{j}^{(\eta_{L})})^{2}]=\kappa^{2}$. We run the experiment with sparsity level $\alpha=0.5$, variance of the entries of the first signal $\kappa^{2}={8\delta}$, and perturbation to each consecutive signal of variance $\sigma_{w}^{2}=400{\delta}$, for varying $\delta$. This ensures that the signal power ${\mathop{\mathbb{E}}}[(\bm{X}_{i})^{\top}{\bm{\beta}}^{(i)}]^{2}$ and the signal difference ${\mathop{\mathbb{E}}}[(\bm{X}_{i})^{\top}{\bm{\beta}}^{(i)}-(\bm{X}_{i+1})^{\top}{\bm{\beta}}^{({i+1})}]^{2}$ for $i\in\{\eta_{\ell}\}_{\ell=1}^{L-1}$ are held constant for varying $\delta$. There are two change points at $n/3$ and $8n/15$. Without perfect knowledge of the magnitude $\sigma_{w}^{2}$ and sparsity level $\alpha$ of the sparse difference vector, AMP assumes $\hat{\sigma}_{w}^{2}=2500\delta$ and $\hat{\alpha}=0.9$ in the experiments in Figure 3b.
@@ Appendix F Computational Cost
For $i\in[n]$, the function $g_{i}^{*t}$ in Proposition 2 only depends on $\bar{{\bm{\Psi}}}$ in (88) through the marginal probability of $\bar{\Psi}_{i}$, that is, $\pi_{\bar{\Psi}_{i}}:\ell\mapsto\sum_{\{{\bm{\psi}}:\psi_{i}=\ell\}}\pi_{\bar{{\bm{\Psi}}}}({\bm{\psi}})$. This means $g_{i}^{*t}$ can be efficiently computed, involving only a sum over $\bar{\Psi}_{i}\in[L]$. Indeed, from the implementation details in Appendix C.2, both $f_{j}^{*t}$ and $g_{i}^{*t}$ can be computed in $O(L^{3})$ time for each $i,j$. Thus $f^{*t}$ and $g^{*t}$ can be computed in $O(nL^{3})$. The per-iteration computational cost of AMP is therefore dominated by the matrix multiplications in (LABEL:eq:amp), which are $O(npL)$.
@@
For ${\bm{\Psi}}\in\mathcal{X}$, the denominator in (33) only requires evaluating a polynomial number of terms, $n\choose{L-1}$}. Therefore, choosing $f^{t},g^{t}$ as in Proposition 2, the computational complexity of computing the quantity $p_{\bar{{\bm{\Psi}}}|\bar{\bm{V}}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}({\bm{\psi}}|\bm{\Theta}^{t},\bm{y})$ from (35) along with AMP is $O(npL+nL^{3}+n^{L-1})=O(np+n^{L-1})$. Further, assuming the expectations in (10)–(13) can be well-approximated in $O(n)$ time, $p_{\bar{{\bm{\Psi}}}|{\bm{V}}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}({\bm{\psi}}|\bm{\Theta}^{t},\bm{y})$ from (33) along with AMP can be computed in $O(np+n^{L-1})$ time.
Appendix G Background on AMP for Generalized Linear Models
Here we review the AMP algorithm and its state evolution characterization for the standard Generalized Linear Model (GLM) without change points. As in Section 3, let $\bm{B}\in{\mathbb{R}}^{p\times L}$ be a signal matrix and let $\bm{X}\in{\mathbb{R}}^{n\times p}$ be a design matrix. The observation $\bm{y}\in{\mathbb{R}}^{n\times L}$ is produced as
$\bm{y}=q(\bm{X}\bm{B},\bm{\varepsilon})\in{\mathbb{R}}^{n\times L},$ (124)
where $\bm{\varepsilon}\in{\mathbb{R}}^{n}$ is a noise vector, and $q:{\mathbb{R}}^{L}\to{\mathbb{R}}^{L}$ is a known output function. The only difference between this model and the one in (5) is the absence of the signal configuration vector $\bm{\Psi}$}. The AMP algorithm for the GLM in (124) was derived by Rangan (2011) for the case of vector signals ($L=1$); see also Section 4 of Feng et al. (2022). Here we discuss the algorithm for the general case ($L\geq 1$), which can be found in Tan and Venkataramanan (2023). For ease of exposition, we will make the following standard assumption (see (S0) on p.(S0)): as $n,p\to\infty$, the empirical distributions of $\{\bm{B}_{j}\}_{j\in[p]}$ and $\{\varepsilon_{i}\}_{i\in[n]}$ converge weakly to laws ${\mathbb{P}}_{\bar{\bm{B}}}$ and ${\mathbb{P}}_{\bar{\varepsilon}}$, respectively, with bounded second moments. We also recall that $n/p\to\delta$ as $n,p\to\infty$.
The AMP algorithm for the model (124) is the same as the one in (LABEL:eq:amp), but due to the assumption above, we can take $f^{t},g^{t}$ to be separable, that is, $f^{t}:{\mathbb{R}}^{L}\to{\mathbb{R}}^{L}$ and $g^{t}:{\mathbb{R}}^{L}\times{\mathbb{R}}^{L}\to{\mathbb{R}}^{L}$ act row-wise on their matrix inputs. Then, the matrices $\bm{F}^{t}$ and $\bm{C}^{t}$ in (LABEL:eq:amp) can be simplified to $\bm{C}^{t}=\frac{1}{n}\sum_{i=1}^{n}\textrm{d}g^{t}(\bm{\Theta}_{i}^{t},y_{i})$ and $\bm{F}^{t}=\frac{1}{n}\sum_{j=1}^{p}\textrm{d}f^{t}(\bm{B}_{j}^{t})$, where $\textrm{d}g^{t}$ and $\textrm{d}f^{t}$ denote the $L\times L$ Jacobians with respect to the first argument.
State evolution
The memory terms $\hat{{\bm{R}}}^{t-1}(\bm{F}^{t})^{\top}$ and $\hat{\bm{B}}^{t}(\bm{C}^{t})^{\top}$ in (LABEL:eq:amp) debias the iterates $\bm{\Theta}^{t}$ and $\bm{B}^{t+1}$ and enable a succinct distributional characterization, guaranteeing that their empirical distributions converge to well-defined limits as $n,p\to\infty$. Specifically, Theorem 1 in Tan and Venkataramanan (2023) shows that for each $t\geq 1$, the empirical distribution of the rows of $\bm{B}^{t}$ converges to the law of a random vector $\tilde{\bm{V}}_{\bm{B}}^{t}:=\bar{\bm{B}}\tilde{{\bm{\nu}}}^{t}_{\bm{B}}+\tilde{\bm{G}}^{t}_{\bm{B}}\,\in{\mathbb{R}}^{1\times L}$, where $\tilde{\bm{G}}^{t+1}_{\bm{B}}\sim{\sf N}(0,\tilde{{\bm{\kappa}}}_{\bm{B}}^{t,t})$ is independent of $\bar{\bm{B}}\sim{\mathbb{P}}_{\bar{\bm{B}}}$. Similarly, recalling that $\bm{\Theta}=\bm{X}\bm{B}\in{\mathbb{R}}^{n\times L}$, the empirical distribution of the rows of $(\bm{\Theta},\,\bm{\Theta}^{t})$ converges to the law of the random vectors $(\tilde{{\bm{Z}}},\tilde{{\bm{Z}}}\tilde{{\bm{\nu}}}^{t}_{\bm{\Theta}}+\tilde{\bm{G}}^{t}_{\bm{\Theta}})$, where $\tilde{\bm{G}}^{t+1}_{\bm{\Theta}}\sim{\sf N}(0,\tilde{{\bm{\kappa}}}_{\bm{\Theta}}^{t,t})$ is independent of $\tilde{{\bm{Z}}}\sim{\sf N}\left(0,\delta({\mathop{\mathbb{E}}}[\bar{{\bm{B}}}\bar{{\bm{B}}}^{\sf T}])^{-1}\right)$. The deterministic $L\times L$ matrices $\tilde{{\bm{\nu}}}^{t}_{\bm{B}},\tilde{{\bm{\kappa}}}_{\bm{B}}^{t,t},\tilde{{\bm{\nu}}}_{\bm{\Theta}}^{t,t}$, and $\tilde{{\bm{\kappa}}}_{\bm{\Theta}}^{t,t}$ can be recursively computed for $t\geq 1$ via a state evolution recursion that depends on $f^{t},g^{t}$ and the limiting laws ${\mathbb{P}}_{\bar{\bm{B}}}$ and ${\mathbb{P}}_{\bar{\varepsilon}}$.
The state evolution characterization allows us to compute asymptotic performance measures such as the MSE of the AMP algorithm. Indeed, for each $t\geq 1$, we almost surely have $\lim_{p\to\infty}\frac{1}{p}\|f^{t}(\bm{B}^{t})-\bm{B}\|_{F}^{2}={\mathop{\mathbb{E}}}[\|f^{t}(\tilde{\bm{V}}_{\bm{B}}^{t})-\bar{\bm{B}}\|^{2}_{2}]$, where the expectation on the right can be computed using the joint law of the $L$-dimensional random vectors $\bar{\bm{B}}$ and $\tilde{\bm{V}}_{\bm{B}}^{t}=\bar{\bm{B}}\tilde{{\bm{\nu}}}^{t}_{\bm{B}}+\tilde{\bm{G}}^{t}_{\bm{B}}$.
In the model (3) with change points, we have an additional signal configuration vector $\bm{\Psi}$}, because of which we cannot take the AMP denoising function $g^{t}$ to be separable, even under Assumption (S0). This leads to a more complicated state evolution characterization, as described in Section 3.